OpenStudy (anonymous):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

stressing

OpenStudy (anonymous):

any 1

OpenStudy (anonymous):

;(

jimthompson5910 (jim_thompson5910):

" Sam rented a boat at $225 for 2 days." so what are x and y?

OpenStudy (anonymous):

guess at 2 days is the y

OpenStudy (anonymous):

and 225 x?

jimthompson5910 (jim_thompson5910):

go back to the definition of what x and y are (they set that up)

jimthompson5910 (jim_thompson5910):

"Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days."

OpenStudy (anonymous):

ahh days x

jimthompson5910 (jim_thompson5910):

that's why it's handy to say x = number of days y = total cost

jimthompson5910 (jim_thompson5910):

" Sam rented a boat at $225 for 2 days." x = ?? y = ??

OpenStudy (anonymous):

2x and 225y

jimthompson5910 (jim_thompson5910):

2x is not correct

jimthompson5910 (jim_thompson5910):

x = 2x makes no sense really

OpenStudy (anonymous):

i dont know what your asking?

OpenStudy (anonymous):

im soo badd at maths :( sorry

jimthompson5910 (jim_thompson5910):

x is simply a number

jimthompson5910 (jim_thompson5910):

x is defined to be the number of days

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so saying x = 2x makes no sense, but you're close

OpenStudy (anonymous):

just 2x?

OpenStudy (anonymous):

something like y =2x=225?

jimthompson5910 (jim_thompson5910):

no don't worry about the equation right now

jimthompson5910 (jim_thompson5910):

we're just setting up the points

jimthompson5910 (jim_thompson5910):

x is the number of days

jimthompson5910 (jim_thompson5910):

if he rents the boat for 2 days, then x = ?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

yep x = 2

jimthompson5910 (jim_thompson5910):

what is y when x = 2

OpenStudy (anonymous):

225

jimthompson5910 (jim_thompson5910):

they pair up together to get the ordered pair (x,y) = (2,225) This is one point

jimthompson5910 (jim_thompson5910):

what is the other point?

OpenStudy (anonymous):

5,48-

OpenStudy (anonymous):

480*

jimthompson5910 (jim_thompson5910):

use parenthesis when you are referring to points or ordered pairs

jimthompson5910 (jim_thompson5910):

so (5,480)

jimthompson5910 (jim_thompson5910):

we have 2 points: (2,225) and (5,480)

jimthompson5910 (jim_thompson5910):

find the equation of the line that goes through them

OpenStudy (anonymous):

is that y2-y1 / x2 - x1?

jimthompson5910 (jim_thompson5910):

that will give you the slope

jimthompson5910 (jim_thompson5910):

and you'll use that to find the equation of the line

OpenStudy (anonymous):

(2,225) and (5,480) 480-225 = 255 / 3 = 85

OpenStudy (anonymous):

so is thay x or y? or neither

jimthompson5910 (jim_thompson5910):

that's the slope

jimthompson5910 (jim_thompson5910):

m = 85 is the slope

OpenStudy (anonymous):

so x

OpenStudy (anonymous):

:o

jimthompson5910 (jim_thompson5910):

no, x is not the slope

OpenStudy (anonymous):

i though x was slope:O

jimthompson5910 (jim_thompson5910):

no the slope is change in y over change in x

OpenStudy (anonymous):

:O what the next step then?

jimthompson5910 (jim_thompson5910):

you have the slope and you have a point on the line

jimthompson5910 (jim_thompson5910):

use them to find the y-intercept

OpenStudy (anonymous):

ehat is the equation i should be using?

jimthompson5910 (jim_thompson5910):

y = mx+b is the general form of a straight line

jimthompson5910 (jim_thompson5910):

m is the slope

jimthompson5910 (jim_thompson5910):

b is the y-intercept

OpenStudy (anonymous):

so we know m only? , sorry im so lost

OpenStudy (anonymous):

y = 85x + ?

jimthompson5910 (jim_thompson5910):

remember, each point we set up is in the form (x,y)

jimthompson5910 (jim_thompson5910):

so pick any point we set up, and plug it in

OpenStudy (anonymous):

225=85(2) + ?

jimthompson5910 (jim_thompson5910):

replace the question mark with b

jimthompson5910 (jim_thompson5910):

so 225 = 85(2) + b now solve for b

OpenStudy (anonymous):

225=85(2) + b or is it y = 85(2) + 225

OpenStudy (anonymous):

b =55

jimthompson5910 (jim_thompson5910):

225=85(2) + b, solve for b

jimthompson5910 (jim_thompson5910):

it's not y = 85(2) + 225

OpenStudy (anonymous):

yes 55?

OpenStudy (anonymous):

b =55?

jimthompson5910 (jim_thompson5910):

good, b = 55

jimthompson5910 (jim_thompson5910):

so y = mx+b turns into y = 85x + 55

jimthompson5910 (jim_thompson5910):

you now have to get that into standard form Ax + By = C to answer part a) make sure A is positive

OpenStudy (anonymous):

now im completely lost

jimthompson5910 (jim_thompson5910):

do you see how I'm getting y = 85x + 55

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its the next part

jimthompson5910 (jim_thompson5910):

so you need to get that equation into the form Ax + By = C A,B,C are constants

OpenStudy (anonymous):

thats the bit confusing me

jimthompson5910 (jim_thompson5910):

for example, 9x + 10y = 12 is in the form Ax + By = C where A = 9, B = 10, C = 12 This is not the answer as it's just an example

OpenStudy (anonymous):

85x +55y = C

OpenStudy (anonymous):

i dunno

jimthompson5910 (jim_thompson5910):

start with y = 85x + 55 and move over the x term to the left side

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

how would you move over the +55? You would do the same for the 85x

OpenStudy (anonymous):

subtract?

jimthompson5910 (jim_thompson5910):

yep subtract 85x from both sides.

OpenStudy (anonymous):

ca u show me how

OpenStudy (anonymous):

like i get theory behingg but now how

jimthompson5910 (jim_thompson5910):

give it a shot and show me what you get

jimthompson5910 (jim_thompson5910):